7(x+1)=4[x-(3-x)]
I am not sure how to simplify the right side of the equation. I just can't remember the distribution.

Answer by palanisamy(496)

7(x+1)=4[x-(3-x)]
7x+7=4[x-3+x]
7x+7=4[2x-3]
7x+7=8x-12
7+12=8x-7x
19=x
x=19